A firm’s total revenue function is given by R = 100 + 100Q - 2Q2 . When Q=13 marginal revenue is ___?

Hint: Answer up to the second decimal.
Answer #1

As it has been given that firm’s total revenue function is given by R = 100 + 100Q - 2Q2 .

When Q=13 marginal revenue

MR=dR/dQ

=0+100(1)-2*2Q

=100-4Q

MR when Q=13

MR=100-4*13

=100-52

=48